# TrajAgent: An LLM-based Agent Framework for Automated Trajectory Modeling via Collaboration of Large and Small Models

### v3

@@ A.5. Case Study
Figure 6. A representative example case of TrajAgent.
A.6. Prompt Examples
{mdframed}
[backgroundcolor=gray!20, skipabove=skipbelow=]
Parameter Optimization
User
Please select proper combination of hyperparameters of model in CONFIG HYPERPARAMETERS.Adjust the selection, the combination of hyperparameters based on the main function of hyperparameters, the characteristics of the input data, the tuning principles, and memory to get a high score. You should solve the task with interleaving Thought, Action and Observation steps.
¡CHARACTERISTICS OF INPUT DATA¿
The input temporal data contains a time dictionary(key is the user ID,the value is a list containing all time points when the user is active in chronological order) , the input user data contains a user dictionary(key is the user ID,the value is a list containing all items that the user interacts with in chronological order).
¡CONFIG HYPERPARAMETERS¿
{loc_emb_size: 500, tim_emb_size: 10, ...}
¡TUNING PRINCIPLES¿
1.Start with a small batch size (32-64) and a small learning rate (0.001-0.01): This will help prevent overshooting and overfitting.
2.Increase batch size and learning rate: If the model is not overfitting, increasing the batch size and learning rate can help improve convergence.
3.Add dropout (0.2-0.5) to prevent overfitting: If the model is overfitting, adding dropout can help regularize the model.
4.Increase embedding size: If the model is not capturing enough information, increasing the embedding size can help improve representational power.
5.Decrease learning rate and increase batch size: If the model is not converging, decreasing the learning rate and increasing the batch size can help improve stability.
¡MEMORY¿:
¡SCRATCHPAD¿:
In Thought step,you should reason how to choose the combination of hyperparameters to get a higher score.Please consider following aspects: 1.Observe the hyperparameters with high scores in MEMORY, to determine the optimal hyperparameters.
2.Use a grid search: Perform a grid search over a range of hyperparameters with high scores to find the optimal combination with higher scores.
3.Increase batch size and learning rate, and use dropout to avoid overfitting.
4. Stop or reverse the adjusting trend if the score is decreasing. According to above aspects,please first learn experiences from MEMORY, then make plan for the action step.Please use the sentence structure ’Firstly… Then… Lastly’.Let’s think step by step.
Thought:
Assistant
Firstly, I will analyze the current hyperparameters in the CONFIG HYPERPARAMETERS. …
Then, I will consider the hidden size, which is currently set at 500. …
Next, the learning rate is currently at 0.001, which is a good starting point. …
Lastly, I will monitor the model’s performance closely….
Now, I will proceed with the action step to adjust the hyperparameters based on this reasoning.
Action:
1. Increase the tim_emb_size from 10 to 20 to capture more temporal information.
2. Keep the loc_size at 500 for now, as it seems adequate.
3. Increase the hidden_size from 500 to 600 to enhance the model’s capacity.
4. Increase the dropout_p from 0.3 to 0.4 to further prevent overfitting.
5. Set the learning_rate to 0.005 to improve convergence speed.
6. Start with a batch size of 32 and monitor the performance closely.
Observation:
After implementing these changes, I will evaluate the model’s performance on a validation set. …
User
In Action step, you should consider the Thought step in SCRATCHPAD, and give a dict:{hypermeter name:hypermeter value}.The hypermeter name should be the same with the raw config hypermeter names in CONFIG HYPERPARAMETERS,and hypermeter values should be the same type as the hypermeter values in CONFIG HYPERPARAMETERS.Please do not add any comments to each value.
Action:
Assistant
{’loc_emb_size’: 500,…}
{mdframed}
[backgroundcolor=gray!20, skipabove=skipbelow=]
Task Understanding
User
The description of each task is in TASK_DESCRIPTION.
Please parse out the task name the sentence aims to address in RAW_INSTRUCTS.
1. The task name should match the key in the DESCRIPTION.
2. Please only output the task name.Do not output other contents.
¡TASK_DESCRIPTION¿
{Next_Location_Prediction: …,Trajectory_User_Linkage:…}
¡RAW_INSTRUCT¿
I want to identify the trajectory…
Assistant
Trajectory_Completion
{mdframed}
[backgroundcolor=gray!20, skipabove=skipbelow=]
Data Augmentation
User
¡TASK¿
Please:
1.select proper augmentation methods and use them in proper order to jointly augment the input temporal data and user data.Please adjust the selection and combination sequence of operators based on MEANING OF OPERATORS, CHARACTERISTICS OF INPUT DATA, and MEMORY to get a high score.
2.select proper combination of hyperparameters of each augmentation method in CONFIG HYPERPARAMETERS.Adjust the selection, the combination of hyperparameters based on the main function of hyperparameters, the characteristics of the input data, and memory to get a high score.
You should solve the task with interleaving Thought, Action and Observation steps.
¡CONFIG HYPERPARAMETERS¿
1.Ti-crop: crop_nums:(int) the size of cropping. default is 2.
crop_ratio:(float) the ratio of cropping. default is 0.
crop_n_times:(int) the number of cropping for each sequence for default setting. default is 2.
crop_time_sort:(str, choice in [maximum, minimum]) choose the candidate subsequence in a descending/ascending order according to its time interval variance. default is ’minimum’. …
2.Ti-insert_unvisited: …
¡MEANING OF OPERATORS¿
1.Ti-crop: For each user, filter out sessions with fewer trajectory points than ti_threshold + ti_crop_n_times. In each session, randomly extract several segments (ti_crop_n_times) of equal length (ti_threshold). Sample data segments according to the rules.
2.Ti-insert_unvisited:…
¡MEMORY¿:
¡SCRATCHPAD¿:
In Thought step,you should reason how to choose the combination of operators and proper combination of hyperparameters of each augmentation method to get a higher score.Please consider following aspects:
1.The meaning of each operator,and the impact of adding or removing operators at a specific location on the score.
2.Common characteristics of operator index lists with higher scores in MEMORY.Avoid using the same index list as MEMORY with scores lower than 0.18111313868613138.
3.The meaning of hyperparameters of each operator you select, and how to adjust the hyperparameters based on the characteristics of input data and hyperparameters with high scores in MEMORY.
4.Use a grid search: Perform a grid search over a range of hyperparameters with high scores to find the optimal combination with higher scores.
7. Stop or reverse the adjusting trend if the score is decreasing.
According to above aspects,please first learn experiences from MEMORY, then make plan for the action step.Please use the sentence structure ’Firstly… Then… Lastly’.Let’s think step by step.
Thought:
assistant
**Thought:**
Firstly, I will analyze the characteristics of the input data, which consists of user activity timestamps and the corresponding items they interacted with. ….
Then, I will consider the meaning of each operator and their potential impact on the score. For instance, using ‘Ti-crop‘ can help in segmenting the data, …
Lastly, I will select hyperparameters for each operator based on their functions and the characteristics of the input data.
**Action:**
1. **Select Augmentation Methods:** - **Ti-crop**:
- ‘crop_nums‘: 3
- ‘crop_ratio‘: 0.1 (to allow for a small portion of the data to be cropped)
- ‘crop_n_times‘: 3
2. **Hyperparameter Configuration:**
- For ‘Ti-crop‘, I will keep ‘crop_n_times‘ at 3 to ensure multiple segments are created….
3. **Evaluate and Adjust:**
- After implementing the above configurations, I will evaluate the score…
**Observation:**
I will monitor the score after each adjustment to ensure that the changes lead to an improvement. If the score decreases, I will revert to the previous settings.
User
¡SCRATCHPAD¿
**Action:**
1. **Select Augmentation Methods:**
- **Ti-crop**: …
In Action step, you should consider the Thought step in SCRATCHPAD, and return a list and a dictionary.
The list should contain the indices of augmentation methods in {1: ’Ti-crop’, 2: ’Ti-insert_unvisited…}.For example, if you want to first use Ti-crop, secondly use Ti-insert_random, thirdly use Ti-mask, then the list should be ’[1,4,8]’.
The dict:….The hypermeter name should be the same with the corresponding operator hypermeter names in CONFIG HYPERPARAMETERS,and hypermeter values should be the same type as the corresponding operator hypermeter values in CONFIG HYPERPARAMETERS.
Please directly output the list and dictionary.For example:…
Action:
Assistant
[1, 3, 5]
{{crop_nums:3,crop_ratio:0,….},3:{…},5:{…}}

### v5

@@ A.5 Limitations and Failure Mode Analysis
To address these issues, we propose two enhancements: (1) a contrastive reflection mechanism that learns from both successful and failed trials to avoid redundant exploration; and (2) a dynamic memory management strategy that prunes low-performing historical actions. Experimental results (Tables 12 and 11) confirm that these strategies effectively mitigate performance degradation at large step/memory sizes.
Figure 6: A representative example case of TrajAgent.
A.6 Additional Experimental Analysis
The optimization effect diminishes as the step increases: While selecting combinations of operators, the model further optimizes the configuration of each operator (e.g., the original configuration file for "inserter" is insert_nums: 1, insert_ratio: 0, insert_time_sort: maximum, percent_no_augment: 0, ti_insert_n_times: 1). In a zero-shot scenario, the model explores optimization strategies based on dataset characteristics and the meaning of the operators, with a probability of converging to a local optimum—i.e., finding a suboptimal combination and deeming the result sufficient, thus stopping the exploration of new combinations and selecting the best operator configuration based on this combination. We compared the llama3-70b used in the paper with other reasoning models, and the results are shown in Table 11 (S2TUL, FSK-London, memory-length=1). We found that models with stronger reasoning capabilities attempt more operator combinations and have a higher probability of finding better combinations in fewer steps. For instance, DeepSeek-v3 outperforms LLaMA-3-70B in step 4. Under the same operator combination and the same number of steps, models with stronger reasoning capabilities yield better optimization results. For instance, DeepSeek-v3 outperforms LLaMA-3-70B in steps 3, 5, 6, 7, 9, 10, 11, and 13.
Improvement solution: Implement reflection similar to contrastive learning between steps, such as further adjusting the parameters of each operator based on effective combinations, to avoid exploring ineffective combinations as much as possible. The improved results are shown in Table 12 (S2TUL, FSK-London, memory-length=1).
The optimization effect diminishes as the memory length increases: memory length refers to the number of action proposals generated in each reasoning step. Increasing it can improve exploration efficiency but also raises the probability of falling into a local optimum. We compared the performance of llama3-70b with that of reasoning models, and the results are shown in Table 9 (S2TUL, FSK-London, step=5). Both models achieved relatively good results at memory_length=2, but as the length increased, they faced the issue of converging to suboptimal solutions. However, models with stronger reasoning capabilities exhibited a stronger tendency to explore other combinations, thus having a higher probability of finding better combinations and escaping the "optimization trap" .
Improvement solution: Periodically optimize and update the memory organization, discard poor combinations, retain good ones, and guide the model to explore new combinations during the reflection phase. The improved results are shown in Table 11 (S2TUL, FSK-London, step=5).
Table 8: Token consumption and time cost for trajectory processing with different models ($\text{traj}=200$). Input and output token counts are reported for two inference settings: step=1 and step=5 (with memory_length=1).
Model token(step=1) token(step=5) time cost
input output input output
LLM-ZS 37,327 90,180 194,358 282,270 3h27min
DutyTTE 1,108 284 7,718 1,823 1h17min
Table 9: Step-wise ACC@5 performance and Operator Combination Order(OCO) traces for multiple large reasoning models.
step 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17
Llama3-70b
ACC@5 56.28 56.28 59.10 68.18 68.18 68.39 68.18 69.04 69.04 68.39 69.04 69.04 87.01 69.05 68.40 68.19 83.12 88.96
OCO [] [] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [2,9,10] [1,3,9] [1,3,9] [1,3,9] [2,9,10] [2,9,10]
DeepSeek-v3
ACC@5 56.28 79.22 59.52 61.26 81.17 68.83 69.26 69.05 83.98 70.35 69.05 69.26 71.00 61.90 60.82 68.83 71.00 66.88
OCO [] [2,9,10] [1,3,9] [1,3,9] [2,9,10] [1,3,9] [1,3,9] [1,3,9] [2,9,10] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [1,4,9] [1,3,9] [1,3,9] [1,3,5,9]
Gemini-2.0-flash-001
ACC@5 56.28 73.16 61.69 70.35 68.83 82.25 69.05 68.40 87.01 65.58 79.87 83.12 71.65 84.63 82.47 71.43 69.05 69.05
OCO [] [2,9,10] [1,3,9] [1,3,9] [1,3,9] [3,9,10] [1,3,9] [1,3,9] [2,9,10] [1,3,9] [2,9,10] [2,9,10] [1,3,9] [2,9,10] [2,9,10] [1,3,9] [1,3,9] [1,3,9]
Table 10: Performance (ACC@5) and Operator Combination Order(OCO) traces under different memory lengths for Llama3-70b and DeepSeek-v3.
memory-length 0 1 2 3 4 5 6 7
Llama3-70b
ACC@5 56.28 68.39 80.95 79.00 55.19 83.33 57.79 62.55
OCO [] [1,3,9] [2,9,10] [2,9,10] [1,2,9] [2,9,10] [1,4,9] [2,6,10]
DeepSeek-v3
ACC@5 56.28 79.22 59.09 79.65 61.47 80.74 59.52 81.60
OCO [] [2,9,10] [1,3,9] [3,6,9] [2,5,9] [2,9,10] [1,4,9] [2,9,10]
Table 11: Raw step-wise performance and Operator Combination Order(OCO) traces for llama3-70b before and after improvement.
step 0 1 2 3 4 5 6 7 8 9 10 11
Pre-improvement
ACC@5 56.28 56.28 59.10 68.18 68.18 68.39 68.18 69.04 69.04 68.39 69.04 69.04
OCO [] [] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [1,3,9]
Post-improvement
ACC@5 56.28 59.09 61.03 68.18 69.04 69.04 60.38 69.04 82.90 80.08 87.01 83.12
OCO [] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [2,9,10] [2,9,10] [2,9,10] [2,9,10]
Table 12: Performance and tool invocation Operator Combination Order(OCO) traces under different memory lengths before and after improvement.
memory-length 0 1 2 3 4 5 6 7
Pre-improvement
ACC@5 56.28 68.39 80.95 79.00 55.19 83.33 57.79 62.55
OCO [] [1,3,9] [2,9,10] [2,9,10] [1,2,9] [2,9,10] [1,4,9] [2,6,10]
Post-improvement
ACC@5 56.28 68.39 69.05 80.95 81.60 84.63 80.52 85.73
OCO [] [1,3,9] [1,3,9] [2,9,10] [2,9,10] [2,9,10] [2,9,10] [3,6,9]
Table 13: Performance comparison of TrajAgent with and without utilizing score feedback in memory across optimization steps.
Step 0 1 2 3 4 5 6 7 8 9 10 11 12
With score in memory
ACC@5 (%) 56.28 56.28 59.10 68.18 68.18 68.39 68.18 69.04 69.04 68.39 69.04 69.04 87.01
Operators [] [] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [1,3,9] [2,9,10]
Without score in memory
ACC@5 (%) 56.28 45.23 78.14 47.40 48.05 57.36 47.61 58.22 58.23 58.87 58.87 46.32 63.85
Operators [] [1,4,9] [2,9,10] [1,4,9] [1,4,9] [1,4,9] [1,4,9] [1,4,9] [1,4,9] [1,4,9] [1,4,9] [1,4,9] [2,6,10]
A.7 Prompt Examples
Parameter Optimization User Please select proper combination of hyperparameters of model in CONFIG HYPERPARAMETERS.Adjust the selection, the combination of hyperparameters based on the main function of hyperparameters, the characteristics of the input data, the tuning principles, and memory to get a high score. You should solve the task with interleaving Thought, Action and Observation steps. <CHARACTERISTICS OF INPUT DATA> The input temporal data contains a time dictionary(key is the user ID,the value is a list containing all time points when the user is active in chronological order) , the input user data contains a user dictionary(key is the user ID,the value is a list containing all items that the user interacts with in chronological order). <CONFIG HYPERPARAMETERS> {loc_emb_size: 500, tim_emb_size: 10, ...} <TUNING PRINCIPLES> 1.Start with a small batch size (32-64) and a small learning rate (0.001-0.01): This will help prevent overshooting and overfitting. 2.Increase batch size and learning rate: If the model is not overfitting, increasing the batch size and learning rate can help improve convergence. 3.Add dropout (0.2-0.5) to prevent overfitting: If the model is overfitting, adding dropout can help regularize the model. 4.Increase embedding size: If the model is not capturing enough information, increasing the embedding size can help improve representational power. 5.Decrease learning rate and increase batch size: If the model is not converging, decreasing the learning rate and increasing the batch size can help improve stability. <MEMORY>: <SCRATCHPAD>: In Thought step,you should reason how to choose the combination of hyperparameters to get a higher score.Please consider following aspects: 1.Observe the hyperparameters with high scores in MEMORY, to determine the optimal hyperparameters. 2.Use a grid search: Perform a grid search over a range of hyperparameters with high scores to find the optimal combination with higher scores. 3.Increase batch size and learning rate, and use dropout to avoid overfitting. 4. Stop or reverse the adjusting trend if the score is decreasing. According to above aspects,please first learn experiences from MEMORY, then make plan for the action step.Please use the sentence structure ’Firstly… Then… Lastly’.Let’s think step by step. Thought: Assistant Firstly, I will analyze the current hyperparameters in the CONFIG HYPERPARAMETERS. … Then, I will consider the hidden size, which is currently set at 500. … Next, the learning rate is currently at 0.001, which is a good starting point. … Lastly, I will monitor the model’s performance closely…. Now, I will proceed with the action step to adjust the hyperparameters based on this reasoning.
Action: 1. Increase the tim_emb_size from 10 to 20 to capture more temporal information. 2. Keep the loc_size at 500 for now, as it seems adequate. 3. Increase the hidden_size from 500 to 600 to enhance the model’s capacity. 4. Increase the dropout_p from 0.3 to 0.4 to further prevent overfitting. 5. Set the learning_rate to 0.005 to improve convergence speed. 6. Start with a batch size of 32 and monitor the performance closely. Observation: After implementing these changes, I will evaluate the model’s performance on a validation set. … User In Action step, you should consider the Thought step in SCRATCHPAD, and give a dict:{hypermeter name:hypermeter value}.The hypermeter name should be the same with the raw config hypermeter names in CONFIG HYPERPARAMETERS,and hypermeter values should be the same type as the hypermeter values in CONFIG HYPERPARAMETERS.Please do not add any comments to each value. Action: Assistant {’loc_emb_size’: 500,…} Task Understanding User The description of each task is in TASK_DESCRIPTION. Please parse out the task name the sentence aims to address in RAW_INSTRUCTS. 1. The task name should match the key in the DESCRIPTION. 2. Please only output the task name.Do not output other contents. <TASK_DESCRIPTION> {Next_Location_Prediction: …,Trajectory_User_Linkage:…} <RAW_INSTRUCT> I want to identify the trajectory… Assistant Trajectory_Completion Data Augmentation User <TASK> Please: 1.select proper augmentation methods and use them in proper order to jointly augment the input temporal data and user data.Please adjust the selection and combination sequence of operators based on MEANING OF OPERATORS, CHARACTERISTICS OF INPUT DATA, and MEMORY to get a high score. 2.select proper combination of hyperparameters of each augmentation method in CONFIG HYPERPARAMETERS.Adjust the selection, the combination of hyperparameters based on the main function of hyperparameters, the characteristics of the input data, and memory to get a high score.
You should solve the task with interleaving Thought, Action and Observation steps. <CONFIG HYPERPARAMETERS> 1.Ti-crop: crop_nums:(int) the size of cropping. default is 2. crop_ratio:(float) the ratio of cropping. default is 0. crop_n_times:(int) the number of cropping for each sequence for default setting. default is 2. crop_time_sort:(str, choice in [maximum, minimum]) choose the candidate subsequence in a descending/ascending order according to its time interval variance. default is ’minimum’. … 2.Ti-insert_unvisited: … <MEANING OF OPERATORS> 1.Ti-crop: For each user, filter out sessions with fewer trajectory points than ti_threshold + ti_crop_n_times. In each session, randomly extract several segments (ti_crop_n_times) of equal length (ti_threshold). Sample data segments according to the rules. 2.Ti-insert_unvisited:… <MEMORY>: <SCRATCHPAD>: In Thought step,you should reason how to choose the combination of operators and proper combination of hyperparameters of each augmentation method to get a higher score.Please consider following aspects: 1.The meaning of each operator,and the impact of adding or removing operators at a specific location on the score. 2.Common characteristics of operator index lists with higher scores in MEMORY.Avoid using the same index list as MEMORY with scores lower than 0.1811. 3.The meaning of hyperparameters of each operator you select, and how to adjust the hyperparameters based on the characteristics of input data and hyperparameters with high scores in MEMORY. 4.Use a grid search: Perform a grid search over a range of hyperparameters with high scores to find the optimal combination with higher scores. 7. Stop or reverse the adjusting trend if the score is decreasing. According to above aspects,please first learn experiences from MEMORY, then make plan for the action step.Please use the sentence structure ’Firstly… Then… Lastly’.Let’s think step by step. Thought: assistant **Thought:** Firstly, I will analyze the characteristics of the input data, which consists of user activity timestamps and the corresponding items they interacted with. This suggests that temporal relationships and user preferences are crucial for effective augmentation. The augmentation methods should enhance these aspects without distorting the underlying patterns. Then, I will consider the meaning of each operator and their potential impact on the score. For instance, using ‘Ti-crop‘ can help in segmenting the data, which is essential for understanding user behavior over time. … Lastly, I will select hyperparameters for each operator based on their functions and the characteristics of the input data. **Action:** 1. **Select Augmentation Methods:** - **Ti-crop**: - ‘crop_nums‘: 3 - ‘crop_ratio‘: 0.1 (to allow for a small portion of the data to be cropped) - ‘crop_n_times‘: 3 2. **Hyperparameter Configuration:** - For ‘Ti-crop‘, I will keep ‘crop_n_times‘ at 3 to ensure multiple segments are created…. 3. **Evaluate and Adjust:** - After implementing the above configurations, I will evaluate the score… **Observation:** I will monitor the score after each adjustment to ensure that the changes lead to an improvement. If the score decreases, I will revert to the previous settings. User <SCRATCHPAD> **Action:** 1. **Select Augmentation Methods:** - **Ti-crop**: … In Action step, you should consider the Thought step in SCRATCHPAD, and return a list and a dictionary. The list should contain the indices of augmentation methods in {1: ’Ti-crop’, 2: ’Ti-insert_unvisited…}.For example, if you want to first use Ti-crop, secondly use Ti-insert_random, thirdly use Ti-mask, then the list should be ’[1,4,8]’. The dict:….The hypermeter name should be the same with the corresponding operator hypermeter names in CONFIG HYPERPARAMETERS,and hypermeter values should be the same type as the corresponding operator hypermeter values in CONFIG HYPERPARAMETERS. Please directly output the list and dictionary.For example:… Action: Assistant [1, 3, 5] {{crop_nums:3,crop_ratio:0,….},3:{…},5:{…}}